I am x old now, my brother is 3 times older. How old was my brother last year?
Answer by Alan3354(69443)

I am x old now, my brother is 3 times older. How old was my brother last year?
------------
1 time older = x+x = 2x
2 times older = x+2x = 3x
3 times older = x+3x = 4x
------
Last year he was 4x-1 years old.